9. A machinist produces 22 items during a shift. Three of the 22 items are defective and the rest are not defective. In how many different orders can the 22 items be arranged if all the defective items are considered identical and all the non-defective items are identical of a different class?
Answer #1

ans) The number of ways of arranging 3 defective items and (22-3) =19 non defective items is 22C3.

22c3= 22 (31) (19!)

=(22)(21)(20)/(3)(2)(1)

=1540